OpenStudy (anonymous):

What are the exact solutions of x^2 = 4 - 7x?

OpenStudy (callisto):

x^2 = 4 - 7x x^2 +7x -4=0 \[x=\frac{-7\pm\sqrt{7^2-4(1)(-4)}}{2(1)} =?\]

OpenStudy (anonymous):

u lost me there bro

OpenStudy (callisto):

._. sister... \[x=\frac{-7 \pm \sqrt{7^2 - 4(1)(-4)}}{2(1)} = \frac{-7 \pm \sqrt{65}}{2} \]Done